The base of a traingle is 1/3 as long as its height. If the area is 24 square meters, how long is its base?
Answer by mathslover(157)

Let the height be x
So the base is x/3 (One third of the height )
Area of a triangle
=1/2 * base * height
According to the given condition
1/2 * x/3 * x = 24
x^2/6 = 24
Multiplying both sides by 6
x^2 = 24*6
x^2=144
x=sqrt(144)
therefore x=12
and x/3 = 4
Base is 4 and height is 12